# American Sign Language Translation ASLP American Sign Language Production

### v2

@@
SignDiff: Diffusion Models for American Sign Language Production
@@
Hongbin Zhong5, Mingyu Zhao1, Yapeng Tian6, Chen Chen7
@@ Abstract
Figure 1: The Process of Sign Language Production: (a) Enter an English text you want to translate, our ASLP method then translates the text into a continuous skeleton pose sequence. (b) Use our FR-Net to predict the generated human body shape. Unlike previous pose2video work, FR-Net learn human body shape as an additional input, rather than just skeletal condition; (c) User can use pose-condition human synthesis model [23] to further generate an final sign language video.
@@ II-B Rendering of Conditional Input
Conditioning refers to the ability of a generative model to control its output according to our intentions like ControlNet [73]. Previous conditional input Generative Adversarial Networks (GANs) [19] demonstrated good performance in image [25, 40, 66, 76, 38] and video generation [35, 61, 63, 64, 65]. There have also been numerous works focusing on generating human poses conditioned on various factors, including whole body [1, 34, 38, 52, 57, 77, 8], face [14, 31, 71, 60, 70, 74] and hand [33, 58, 69].
@@ III-B Dataset Processing
The How2Sign dataset is the first large-scale multimodal and multiview continuous ASL dataset. It includes a parallel corpus comprising over 80 hours of sign language videos and corresponding modalities such as speech, English transcripts, and depth information. Despite its significance, prior researchers in the field of sign language have not effectively employed this corpus due to the intricacies involved in American Sign Language Production (ASLP). Our objective is to train a Sign Language Pose network to generate sequences of 3D skeleton poses on How2Sign as shown in Fig 2. Initially, we extract 2D joint positions from each video using the OpenPose method [7]. We then incorporate the skeletal model estimation improvements introduced in Zelink et al. [72] to transform the 2D joint positions into their corresponding 3D counterparts. Moreover, we employ an iterative inverse kinematics approach to minimize the 3D pose while preserving consistent bone lengths and correcting any misplaced joints. Finally, following a similar approach proposed in Stoll et al. [54], we normalize the skeletal data and adjust the data to a trainable state. Our scripts and preprocessing tools are available in supplementary materials and will be useful to the sign language community.
@@ III-C Preliminary
Figure 3: We condition Stable Diffusion model (SD) via concatenation or by a more general cross-attention mechanism, which is now the main theoretical basis for controlling SD.
@@
III-D Tensor-Optimization-Based New Text2Pose Method
@@
Scale-Adjusted Loss. It aims to improve the generalization of our model and address the overfitting problem when handling extremely large-scale datasets, which is then calculated using a specific criterion, which takes the $preds_{masked}$ and $targets_{masked}$ mentioned in the previous paragraph as inputs. To adjust the magnitude of the loss, a loss scale factor is applied. If the loss scale is not equal to 1.0, the loss is multiplied by the loss scale factor. Additionally, SA regularization is performed to prevent overfitting. The regularization loss, $SA_{loss}$, is initialized as a tensor with a value. The SA_loss is initialized only before the optimization loop begins. However, SA_loss is updated within the loop by adding the SA norm of each weight parameter containing the string ’weight’. The SA norm of the weight parameters in the model is calculated using the torch.norm function. The computed SA norm is then added to the $SA_{loss}$ variable for each weight parameter that contains the ‘weight’ string in its name. A regularization coefficient, $SA_{lambda}$, is defined with a value of 0.01 to control the strength of the regularization. The SA regularization loss, the result of $SA_{lambda}$ $SA_{loss}$ will be added to the total loss.
III-E SignDiff for Efficient Pose2Video Production
@@
FR-Net utilizes adaptive thresholding to create a binary image based on local pixel thresholds. Subsequently, a 5x5 convolutional kernel is utilized for erosion, shrinking white regions in the binary image. This helps refine foreground object segmentation. ControlNet [73] locks all parameters in $\Theta$ and then clones it into a trainable copy $\Theta_{\text{c}}$. The copied $\Theta_{\text{c}}$ is trained with an external condition vector $\bm{c}$. We extend a copy of $\Theta_{\text{d}}$ and define the new input as $\bm{d}$. Here $\bm{c}$ represents OpenPose’s input to the model, and $\bm{d}$ represents our FR-Net’s input to the model. Since our input is ultimately superimposed, we can call the variable that $\bm{c}$ and $\bm{d}$ superimpose $\bm{e}$. Overall, we follow the same principles as ControlNet to incorporate new input conditions.
@@
TABLE II: Fast-SLP Transformer results for Text to Sign Pose production on the How2Sign, with multiple data augmentation techniques. FP: Future Prediction, GN: Gaussian Noise, BO: Batch Optimization, CO: Computational Optimization.
@@
Future Prediction 12.67 17.33 25.93 43.68 45.14 12.82 15.05 23.88 41.36 43.24
Just Counter 14.05 20.16 27.71 45.35 47.20 12.80 15.72 25.36 43.76 44.96
Gaussian Noise 15.12 21.92 29.98 47.32 49.74 14.57 18.86 27.03 44.17 46.30
@@
Scale-Adjusted Loss 14.42 20.84 28.33 44.77 47.08 15.12 18.66 26.51 44.03 44.24
BO & CO 16.92 22.43 30.01 45.15 48.75 12.28 15.99 22.67 37.58 37.84
IV Experiments
@@ IV-A Experimental Setup
Our translation model utilizes the video-level translation model developed by Tarrés et al. [59] and translation model [9], which are refined with retraining. However, due to their model’s lower performance than previous best level of mature sign language interpretation, it fails to fully showcase the accuracy of our modelf in ASL tasks. Thus, we hope that future researchers can further explore and expand in American Sign Language Translation (ASLT).
IV-B Evaluation for Fast-SLP
@@
In Table III, we also used previous method to train and test it on How2Sign ASL datasets for comparison with previous work. We now use one of these jobs as a baseline to run our new data set for comparison. The result shows that our approach is also competitive for excellent existing work, multiple sign languages, not just for data quality reasons. NSA [2], which appears after our ASLP work, shows that their performance is only roughly on par with ours.
@@
(a) Inputs (b) Generated images
Figure 6: We presented a comparison of SignDiff with previous work [73, 65], including (a) input and (b) generated images. We demonstrated the effect of training 10,000 images and 50,000 images, which demonstrates that we only need fewer samples and have higher training efficiency. We also demonstrated the effects that two models of native stable diffusion can achieve for comparison. In reality style, sign language is accurate but prone to missing fingers, while in anime style, fingers are complete but not accurate.
@@
The most recent two approaches, Photo-realistic SLP and our proposed Fast-SLP Transformers, showcase the latest advancements in the field. Photo-realistic SLP achieved a notable performance leap on the test set, particularly in terms of BLEU-4 scores. However, our Fast-SLP Transformers method outperformed all previous work across all metrics, achieving the highest BLEU-4 and ROUGE scores on both the development and test sets. This indicates that our approach has made substantial progress in generating more accurate and fluent sign language translations, setting a new benchmark for future research in this area.
IV-C Evaluation for SignDiff
@@
Qualitative Evaluations. Figure 6 showcases three distinct input types and the corresponding generation results from different models. A comparison between our model and Vid2Vid reveals that our approach achieves superior results with less data input. Upon convergence, our model surpasses the performance of Vid2Vid in terms of generation quality. Moreover, when contrasted with a common LoRA [23] model used to generate human figures, our approach demonstrates clear advantages. While stable diffusion models excel in producing accurate gestures, they often lack detail in finger representation. Conversely, the cartoon style preserves finger details but sacrifices some precision in overall gesture depiction. Our model successfully combines the strengths of both approaches, offering both gesture accuracy and exceptional image clarity. Furthermore, as discussed in Section III-A, our method demonstrates superior image quality upper bounds and enhanced stability. This comprehensive improvement in various aspects of image generation underscores the effectiveness of our approach in addressing the challenges of sign language synthesis.
@@
V CONCLUSION

### v3

@@
SignDiff: Diffusion Model for American Sign Language Production
@@
Hongbin Zhong5, Yapeng Tian6, Chen Chen7
@@ Abstract
Figure 1: Our Sign Language Production: First, input the English text you want to translate, and our new Text2Pose method will translate the text into continuous skeletal pose sequences. Then we plot the pose data of each frame into appropriate images, and users can further generate the final video using our new pose-conditioned human synthesis model (SignDiff).
Collaborator Author.
@@ II-B Rendering of Conditional Input
Conditioning refers to the ability to control its output according to our intentions like ControlNet [71]. Previous conditional input Generative Adversarial Networks (GANs) [19] showed good performance in image [24, 39, 64, 74, 37] and video generation [34, 59, 61, 62, 63]. There have also been numerous works focusing on generating human poses conditioned on various factors, including whole body [1, 33, 37, 51, 56, 75, 8], face [14, 30, 69, 58, 68, 72] and hand [32, 57, 67].
@@ III-B Dataset Processing
The How2Sign dataset is the first large-scale multimodal and multiview continuous ASL dataset. It includes a parallel corpus comprising over 80 hours of sign language videos and corresponding modalities such as speech, English transcripts, and depth information. Despite its significance, prior researchers in the field of sign language have not effectively employed this corpus due to the intricacies involved in American Sign Language Production (ASLP). Our objective is to train a Sign Language Pose network to generate sequences of 3D skeleton poses on How2Sign as shown in Fig 2. Initially, we extract 2D joint positions from each video using the OpenPose method [7]. We then incorporate the skeletal model estimation improvements introduced in Zelink et al. [70] to transform the 2D joint positions into their corresponding 3D counterparts. Moreover, we employ an iterative inverse kinematics approach to minimize the 3D pose while preserving consistent bone lengths and correcting any misplaced joints. Finally, following a similar approach proposed in Stoll et al. [53], we normalize the skeletal data and adjust the data to a trainable state.
@@ III-C Preliminary
Figure 3: We condition Stable Diffusion (SD) via concatenation or by a more general cross-attention mechanism, which is now the main theoretical basis for controlling SD. $e_{j}$ is the combined skeleton pose and predicted shape representation, how to get it will be covered in detail in Sec. IV-B.
@@
IV IMPLEMENTATION DETAILS
IV-A Tensor-Optimization-Based New Text2Pose Method
@@
Scale-Adjusted Loss. It aims to improve the generalization of our model and address the overfitting problem when handling extremely large-scale datasets, which is then calculated using a specific criterion, which takes the $preds_{masked}$ and $targets_{masked}$ mentioned in the previous paragraph as inputs. To adjust the magnitude of the loss, a loss scale factor is applied. If the loss scale is not equal to 1.0, the loss is multiplied by the loss scale factor. Additionally, $SA_{loss}$ use regularization to prevent overfitting. The regularization loss, $SA_{loss}$, is initialized as a tensor with a value. The $SA_{loss}$ is initialized only before the optimization loop begins. However, $SA_{loss}$ is updated within the loop by adding the SA norm of each weight parameter containing the string ’weight’. The SA norm of the weight parameters in the model is calculated using the torch.norm function. The computed SA norm is then added to the $SA_{loss}$ variable for each weight parameter that contains the ‘weight’ string in its name. A regularization coefficient, $SA_{lambda}$, is defined with a value of 0.01 to control the strength of the regularization. The SA regularization loss, the result of $SA_{lambda}$ $SA_{loss}$ will be added to the total loss.
IV-B SignDiff for Efficient Pose2Video Production
@@
FR-Net utilizes adaptive thresholding to create a binary image based on local pixel thresholds. Subsequently, a 5x5 convolutional kernel is utilized for erosion, shrinking white regions in the binary image. This helps refine foreground object segmentation. ControlNet [71] locks all parameters in $\Theta$ and then clones it into a trainable copy $\Theta_{\text{c}}$. The copied $\Theta_{\text{c}}$ is trained with an external condition vector $\bm{c}$. We extend a copy of $\Theta_{\text{d}}$ and define the new input as $\bm{d}$. Here $\bm{c}$ represents OpenPose’s input to the model, and $\bm{d}$ represents our FR-Net’s input to the model. Since our input is ultimately superimposed, we can call the variable that $\bm{c}$ and $\bm{d}$ superimpose $\bm{e}$, as shown in Fig. 3. Overall, we follow the same principles as ControlNet to incorporate new input conditions.
@@
TABLE II: Fast-SLP Transformer results for Text to Sign Pose production on the How2Sign, with multiple data augmentation techniques. Lines 2 show the techniques from previous work [44], and the last two lines are our enhanced methods. FP: Future Prediction, GN: Gaussian Noise, BO: Batch Optimization, CO: Computational Optimization.
@@
Scale-Adjusted Loss (Ours) 14.42 20.84 28.33 44.77 47.08 15.12 18.66 26.51 44.03 44.24
BO & CO (Ours) 16.92 22.43 30.01 45.15 48.75 12.28 15.99 22.67 37.58 37.84
V EXPERIMENTS
@@ V-A Experimental Setup
V-B Evaluation for Fast-SLP
@@
In Table III, we also used previous method to train and test it on How2Sign ASL datasets for comparison with previous work. We now use one of these jobs as a baseline to run our new data set for comparison. It shows that our approach is also competitive for excellent latest work, multiple sign languages, not just for data quality reasons. NSA [2], which appears after our ASLP work, shows that their performance is only roughly on par with ours (it just 0.27 more).
@@
Figure 6: Our work is divided into two parts: text2pose and pose2video. In this figure, we provide qualitative images from Fast-SLP Transformers and SignDiff respectively. (a) shows a comparison between predicted pose images and actual pose images, (b) compares our SignDiff rendered images with the original frames. SignDiff uses 2D skeleton visualization images as input to ensure fairness, allowing us to focus on comparing our reconstruction quality.
@@
The most recent two approaches, Photo-realistic SLP and our proposed Fast-SLP Transformers, showcase the latest advancements in the field. Photo-realistic SLP achieved a notable performance leap on the test set, particularly in terms of BLEU-4 scores. However, our Fast-SLP Transformers method outperformed all previous work in all metrics, achieving the highest BLEU-4 and ROUGE scores in both development and test sets. The performance of our text2pose model can be directly observed Figure 6 (a), which indicates that our approach has made substantial progress in generating more accurate and fluent sign language translations, setting a new benchmark for future research in this area.
V-C Evaluation for SignDiff
@@
Qualitative Evaluations. Figure 6 (b) shows the quality comparison between our SignDiff reconstructed videos and the original videos. To ensure fairness, we used the same 2D pose visualization images as input, and our reconstruction quality is almost identical to the original. Furthermore, the diffusion model does not produce motion blur, with each frame showing very clear finger details. Our only limitation, as shown in the sixth column, is that hand orientations can become confused, but this minor flaw does not overshadow the overall achievements.
@@
VI DISCUSSION
We note a concurrent work [2] published before us that adopts an SMPL-based technical approach rather than skeletal poses. Both approaches have their advantages and disadvantages. While the former may have advantages in pure motion generation, for sign language which requires more user-centric approaches, skeletal poses currently can render more realistic and appealing videos for users.
VII CONCLUSION